OpenStudy (anonymous):

what is the number of CO2 molecules in 1.33 moles of CO2

OpenStudy (chmvijay):

1 mole contains 6.023*10^23 molecules of CO2 1.33 moles contain =1.33*6.023*10^23 = 8*10^23 molecules of CO2